# S-Duality for Non-Abelian Monopoles

### v1

@@ 4 Quantum mechanics on $\mathcal{M}(m,\Phi_{0})$
Solutions of (4.2) are specified by $\Phi_{0}$ in the boundary condition (2.3) and the topological charge $m=(m_{1},\ldots,m_{t})$. For the fixed $\Phi_{0}$ and $m$, all solutions (modulo local gauge transformations) form a connected space $\mathcal{M}(m,\Phi_{0})$. As reviewed in Section 2,
@@
In our setting, $\mathcal{M}(m,\Phi_{0})$ is not necessarily hyper-Kähler, so the above argument does not apply. Nevertheless, the same index computation in [21, 22] that yields $\text{dim}\mathcal{M}(m,\Phi_{0})$ for the bosonic moduli also fixes the number of fermionic zero modes [43, 44] (including the non-normalizable modes). In $\mathcal{N}=4$ SYM this number is always $2\,\text{dim}\mathcal{M}(m,\Phi_{0})$, even when $\text{dim}\mathcal{M}(m,\Phi_{0})$ is not divisible by four. We will therefore continue to use (4.31) as our fermionic zero-mode basis.
@@
where $\nabla_{t}\xi_{\alpha}^{n}=\dot{\xi}_{\alpha}^{n}+\Gamma^{n}{}_{pl}\,\dot{z}^{p}\xi_{\alpha}^{l}$. (4.35) is the standard Lagrangian of a one-dimensional supersymmetric sigma model with at least two real supercharges [42]. For completeness, we derive (4.35) via a collective coordinate expansion in Appendix C, without assuming any hyper-Kähler structure [^7] As in $\mathcal{N}=2$ gauge theories [6], integrating out the additional five scalars can contribute to the effective action even when they do not acquire vacuum expectation values. A collective-coordinate expansion in $\mathcal{N}=4$ SYM that includes these scalars was carried out in [56]; however, that derivation uses the hyper-Kähler structure of the moduli space, which is not available in our case..
@@
In what follows, we construct states and operators on the relative moduli space $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})\equiv M$ that will be used in Section 5. From operators $\hat{z}^{\mu}$ and $\hat{\lambda}_{\mu}$, the position eigenstate $|z\rangle$ and the fermionic vacuum $|\Omega\rangle$ are defined by
@@
$O[f]|f^{\prime}\rangle=|f\wedge f^{\prime}\rangle,\qquad O^{\dagger}[f]|f^{\prime}\rangle=|(-1)^{(p+1)p^{\prime}}\ast(f\wedge\ast f^{\prime})\rangle=|f\mathbin{\lrcorner}f^{\prime}\rangle,$ (4.55)
@@
Ground states are harmonic forms. Let $\{f_{k}\}^{K}_{k=1}$ be an orthogonal basis of harmonic forms on $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})$. For two basis elements $f_{k}$ and $f_{k^{\prime}}$ of the same degree,
$(\ast f_{k})\wedge f_{k^{\prime}}=(f_{k},f_{k^{\prime}})\,\varepsilon,$ (4.57)
@@
is the Riemannian volume form. If the harmonic forms are all parallel, i.e. $\nabla f_{k}=0$, then $(f_{k},f_{k^{\prime}})$ is a constant and orthogonality yields
@@
Using the Weitzenböck identity
$\Delta=\nabla^{\ast}\nabla+R_{p},$ (4.60)
nonnegativity of the curvature term $(R_{p}f,f)\geq 0$ forces any harmonic $p$-form to be parallel, i.e. $\nabla f=0$ [57]. This applies when $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})=C(k_{1})$ is a symmetric space. In the stratified case of Subsection 3.2, we will likewise show that $(R_{p}f,f)\geq 0$ on each stratum and hence on $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})$ as a whole, so harmonic forms are parallel there as well. (4.59) is valid in the present context.
@@ 5 Generators of magnetic gauge transformations and the $H^{s}\times(H^{\vee})^{s}$ symmetry
As shown in Subsection 3.3, for $\mathcal{N}=4$ SYM with simple gauge group $G$, when $G$ is broken to $H$ with semisimple factor $H^{s}$, the degeneracy of ground states on $\mathcal{M}_{\mathrm{rel}}(q,\Phi_{0})$ equals the dimension of the charge-$q$ W-boson multiplet in the dual theory. To promote this numerical match into a one-to-one correspondence, and moreover, to define operators generating the $(H^{\vee})^{s}$-action, we still need to assign an $(H^{\vee})^{s}$-weight to each harmonic form.
Let us recall some standard facts about equal-rank homogeneous spaces [51, 49, 58]. Let $\mathcal{G}$ be a compact, connected Lie group and $\mathcal{H}\subset\mathcal{G}$ a connected, closed subgroup of the same rank. Then the de Rham cohomology of $\mathcal{G}/\mathcal{H}$ admits a basis of $\mathcal{G}$-invariant harmonic forms $\{\omega_{w}\}$ indexed by Weyl-group cosets:
$w\in W_{\mathcal{G}}/W_{\mathcal{H}}\;\;\longleftrightarrow\;\;\omega_{w}\in\{\text{a harmonic form basis of}\;\;H^{\ast}(\mathcal{G}/\mathcal{H})\}.$ (5.1)
Here $W_{\mathcal{G}}$ and $W_{\mathcal{H}}$ are Weyl groups of $\mathcal{G}$ and $\mathcal{H}$, and each $w$ is chosen as the minimal-length representative in its coset. The degree of $\omega_{w}$ is
@@
with $\ell(w)$ the Coxeter length. In particular, the identity $e\in W_{\mathcal{G}}/W_{\mathcal{H}}$ corresponds to the 0-form $\omega_{e}=1$. The dimension of the cohomology equals the Weyl-group index:
$\text{dim}H^{\ast}(\mathcal{G}/\mathcal{H})=\chi[\mathcal{G}/\mathcal{H}]=|W_{\mathcal{G}}/W_{\mathcal{H}}|=|W_{\mathcal{G}}|/|W_{\mathcal{H}}|.$ (5.3)
@@
$w\in W_{\mathcal{G}}/W_{Z_{\mathcal{G}}(\mathbf{k})}\;\;\longleftrightarrow\;\;\text{wt}_{w}=w(\mathbf{k})\in O(\mathbf{k}),$ (5.4)
@@
$\Bigl\{\;\Omega_{p}\wedge\omega_{w_{p}}\;\Big|\;w_{p}\in W_{H^{s}}/W_{Z_{H^{s}}(\overline{\mathbf{k}}_{i_{p}})},\;p=1,\ldots,N\Bigr\},$ (5.8)
realizing the $(H^{\vee})^{s}$-representation $V_{\Lambda_{q}}$.
@@
$\text{wt}(\Omega_{p}\wedge\omega_{w_{p}})=\text{wt}(\omega_{w_{p}})=w_{p}(\overline{\mathbf{k}}_{i_{p}})\in O(\overline{\mathbf{k}}_{i_{p}}).$ (5.10)
@@
$\deg(\Omega_{p})=\begin{cases}\text{dim}\mathcal{M}_{\mathrm{rel}}(q,\Phi_{0},k_{i}),&q=1,\\[6.0pt] \frac{1}{2}\,\text{dim}\mathcal{M}_{\mathrm{rel}}(q,\Phi_{0},k_{i}),&q>1,\end{cases}$ (5.11)
since for $q=1$ the fiber contributes a top form, whereas for $q>1$ the bound states on $\mathcal{M}_{\mathrm{rel}}(q,\Phi_{0},k_{i})$ are (anti-)self-dual middle-degree harmonic forms [4].
@@
In what follows, for the examples in Subsections 3.1 and 3.2, we explicitly compute the weights of the harmonic forms in the associated $(H^{\vee})^{s}$-representation and construct the magnetic generators of the $(H^{\vee})^{s}$-action on the monopole Hilbert space.
@@ 5.1 $(H^{\vee})^{s}$-weight decomposition of harmonic forms
    $\overline{\mathbf{k}}_{1}=\bm{e}_{N}-\frac{1}{N}\sum^{N}_{a=1},\bm{e}_{a},$ (5.13)
@@
    Let $\omega$ be the Kähler form on $\mathbb{CP}^{N-1}$. An orthogonal basis of harmonic forms is given by
@@
    which is minuscule in $SO(2N+1)$, and is the lowest weight of the $\mathbf{2N}$-representation in $(H^{\vee})^{s}=USp(2N)$.
@@
    An orthogonal basis of harmonic forms can be chosen as
@@
    $w_{-k}(\overline{\mathbf{k}}_{1})=-\bm{e}_{k},\qquad w_{+k}(\overline{\mathbf{k}}_{1})=+\bm{e}_{k},$ (5.30)
@@
    $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})\cong SU(N+1)/S[U(N-1)\times U(2)],\qquad\text{dim}_{\mathbb{R}}\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})=4N-4,$ (5.35)
@@
    An orthogonal basis of harmonic forms is
    $\left\{f_{ij}=\omega^{j-i-1}\wedge\Omega^{i-1}\;\middle|\;1\leq i<j\leq N+1\right\},$ (5.37)
    where $\omega$ is the Kähler form and $\Omega$ is a primitive harmonic 44-form. Minimal-length representatives of $W_{H^{s}}/W_{Z_{H^{s}}(\overline{\mathbf{k}}_{1})}$ can be chosen as
@@
    An orthogonal basis of harmonic forms may be chosen as
    $\{1,\omega,\ldots,\omega^{N-1},\psi,\omega^{N},\omega^{N+1},\ldots,\omega^{2N-2}\},$ (5.46)
    which are denoted by $f_{m}$ for $1\leq m\leq 2N$. Here $\omega$ is the Kähler form and $\psi$ is the unique primitive middle-degree class satisfying $\omega\wedge\psi=0$. The associated minimal-length representatives in $W_{H^{s}}/W_{Z_{H^{s}}(\overline{\mathbf{k}}_{1})}$ are
@@
    $w_{-k}(\overline{\mathbf{k}}_{1})=-\bm{e}_{k},\qquad w_{+k}(\overline{\mathbf{k}}_{1})=+\bm{e}_{k},$ (5.50)
    the weights carried by the harmonic forms in the $SO(2N)$ vector representation are given by
@@
    is the lowest weight of $(\mathbf{2N{+}1})$-representation of $(H^{\vee})^{s}=SO(2N+1)$, but is not minuscule in $USp(2N)$. There is a second solution
@@
    On both $\mathbb{CP}^{2N-1}$ and $\mathbb{C}^{2N}$, $(R_{p}f,f)\geq 0$, so $(R_{p}f,f)\geq 0$ holds everywhere on $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})$ and every harmonic form is parallel by the same Weitzenböck argument after (4.60). The Euler characteristic is
@@
    Since $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})$ is homeomorphic to $\mathbb{CP}^{2N}$, they have the same cohomology ring. The basis of harmonic forms on $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})$ can be taken as [^8] Here we take $\mathbb{CP}^{2N-1}\cup\mathbb{C}^{2N}$ as $\mathbb{C}^{2N}$ with the sphere $\mathbb{S}_{\infty}^{4N-1}$ at infinity quotiented by the Hopf $\mathbb{S}^{1}$-action so that $\mathbb{CP}^{2N-1}$ arises at infinity. Any constant $k$-form on $\mathbb{C}^{2N}$ is parallel and thus harmonic. A form survives the Hopf reduction and descends to a form on $\mathbb{CP}^{2N-1}\cup\mathbb{C}^{2N}$ if and only if its restriction to $\mathbb{S}_{\infty}^{4N-1}$ is basic for the Hopf fibration, i.e. $\mathbb{S}^{1}$-invariant and horizontal. This condition singles out (5.57), where $\tilde{\omega}=\frac{i}{2}\sum_{a=1}^{2N}dz^{a}\wedge d\bar{z}^{a}$ is the flat Kähler form on $\mathbb{C}^{2N}$.
@@
    Let $f_{k}=\tilde{\omega}^{k-1}$. The simple-root generators of $SO(2N+1)$ in this basis are
@@
    As shown in (2), $C(k_{1})\cong SU(N+1)/S[U(N)\times U(1)]$, $C(k_{2})\cong SU(N+1)/S[U(N-1)\times U(2)]$,
@@
    where $\mathrm{Tot}(S^{\vee})$ is locally isomorphic to $\mathbb{C}^{2}\times\mathrm{Gr}(2,N+1)$. $\text{dim}_{\mathbb{R}}\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})=4N$. On each stratum $\mathbb{CP}^{N}$ and $\mathrm{Tot}(S^{\vee})$, $(R_{p}f,f)\geq 0$. So $(R_{p}f,f)\geq 0$ on $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})$ and all harmonic forms are parallel.
@@
    $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})$ and $\mathrm{Gr}(2,N+2)$ have isomorphic cohomology rings. From (5.37) with $N\rightarrow N+1$, a basis of harmonic forms on $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})$ can be chosen as
    $f^{\prime}_{ij}=\tilde{\omega}^{j-i-1}\wedge\tilde{\Omega}^{i-1},\qquad 1\leq i<j\leq N+2,$ (5.70)
    where $\tilde{\omega}$ and $\tilde{\Omega}$ are the harmonic $2$-form and $4$-form seeds. After the relabeling $f_{ij}=f^{\prime}_{i\;j+1}$, we obtain
    $\left\{f_{ij}=\tilde{\omega}^{j-i}\wedge\tilde{\Omega}^{i-1}\;\middle|\;1\leq i\leq j\leq N+1\right\},$ (5.71)

### v2

@@ 4 Quantum mechanics on $\mathcal{M}(m,\Phi_{0})$
Solutions of (4.2) are specified by $\Phi_{0}$ in the boundary condition (2.2) and the topological charge $m=(m_{1},\ldots,m_{t})$. For the fixed $\Phi_{0}$ and $m$, under the framing (2.3), all solutions (modulo local gauge transformations) form a connected space $\mathcal{M}(m,\Phi_{0})$. As reviewed in Section 2,
@@
In our setting, $\mathcal{M}(m,\Phi_{0})$ is not necessarily hyper-Kähler, so the above argument does not apply. Nevertheless, the same index computation in [21, 22] that yields $\text{dim}\mathcal{M}(m,\Phi_{0})$ for the bosonic moduli also fixes the number of fermionic zero modes [43, 44] (including the non-normalizable modes). In $\mathcal{N}=4$ SYM, this number is always $2\,\text{dim}\mathcal{M}(m,\Phi_{0})$, even when $\text{dim}\mathcal{M}(m,\Phi_{0})$ is not divisible by four. We will therefore continue to use (4.31) as our fermionic zero-mode basis.
@@
where $\nabla_{t}\xi_{\alpha}^{n}=\dot{\xi}_{\alpha}^{n}+\Gamma^{n}{}_{pl}\,\dot{z}^{p}\xi_{\alpha}^{l}$. (4.35) is the standard Lagrangian of a one-dimensional supersymmetric sigma model with at least two real supercharges [42]. For completeness, we derive (4.35) via a collective coordinate expansion in Appendix C, without assuming the hyper-Kähler structure [^8] As in $\mathcal{N}=2$ gauge theories [6], integrating out the additional five scalars can contribute to the effective action even when they do not acquire vacuum expectation values. A collective-coordinate expansion in $\mathcal{N}=4$ SYM that includes these scalars was carried out in [56]; however, that derivation uses the hyper-Kähler structure of the moduli space, which is not available in our case..
@@
In the following, we construct states and operators on the relative moduli space $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})\equiv M$ that will be used in Section 5. From operators $\hat{z}^{\mu}$ and $\hat{\lambda}_{\mu}$, the position eigenstate $|z\rangle$ and the fermionic vacuum $|\Omega\rangle$ are defined by
@@
$O[f]{|{f^{\prime}}\rangle}={|{f\wedge f^{\prime}}\rangle},\qquad O^{\dagger}[f]{|{f^{\prime}}\rangle}=\begin{cases}0,&p>p^{\prime},\\[4.0pt] {|{\,f\;\lrcorner\;f^{\prime}\,}\rangle},&p\leq p^{\prime},\end{cases}$ (4.55)
@@
Ground states are harmonic forms. Let $\{f_{k}\}_{k=1}^{K}$ be an orthogonal basis of harmonic forms on $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})$, all of which have even degree. For two basis elements $f_{k}$ and $f_{k^{\prime}}$ of the same degree,
$f_{k}\wedge(\ast f_{k^{\prime}})=(f_{k},f_{k^{\prime}})\,\varepsilon,$ (4.57)
@@
is the Riemannian volume form. If all harmonic forms are parallel with $\nabla f_{k}=0$, [^9] In the single-stratum case, $\mathcal{M}_{\rm rel}(1,\Phi_{0})=C(k_{1})$ is a compact symmetric space, where all harmonic forms are parallel. But harmonic forms on $\mathcal{M}_{\rm rel}(m,\Phi_{0})$ may not be parallel in general. $(f_{k},f_{k^{\prime}})$ is a constant and orthogonality yields
@@ 5 Generators of magnetic gauge transformations and the $H^{s}\times(H^{\vee})^{s}$ symmetry
As shown in Subsection 3.3, for $\mathcal{N}=4$ SYM with simple gauge group $G$, when $G$ is broken to $H$ with semisimple factor $H^{s}$, the degeneracy of ground states on $\mathcal{M}_{\mathrm{rel}}(q,\Phi_{0})$ equals the dimension of the charge-$q$ W-boson multiplet in the dual theory. To promote this numerical match into a one-to-one correspondence, and moreover, to construct operators generating the $(H^{\vee})^{s}$-action, we still need to assign an $(H^{\vee})^{s}$-weight to each harmonic form.
Let us recall some standard facts about equal-rank homogeneous spaces [51, 49, 57, 58]. Let $\mathcal{G}$ be a compact, connected Lie group and $\mathcal{H}\subset\mathcal{G}$ a connected, closed subgroup of the same rank. Then $H^{\ast}(\mathcal{G}/\mathcal{H};\mathbb{R})$ is concentrated in even degrees and
$\sum_{k}\text{dim}H^{k}(\mathcal{G}/\mathcal{H};\mathbb{R})=\chi[\mathcal{G}/\mathcal{H}]=|W_{\mathcal{G}}/W_{\mathcal{H}}|=|W_{\mathcal{G}}|/|W_{\mathcal{H}}|,$ (5.1)
where $W_{\mathcal{G}}$ and $W_{\mathcal{H}}$ are the Weyl groups of $\mathcal{G}$ and $\mathcal{H}$. If moreover $\mathcal{H}=Z_{\mathcal{G}}(S)$ is the centralizer of a torus $S\subset\mathcal{G}$, then $\mathcal{G}/\mathcal{H}$ is a generalized flag manifold. In this case the Schubert classes form a basis of $H^{\ast}(\mathcal{G}/\mathcal{H};\mathbb{R})$ indexed by the minimal-length representatives $w$ of $W_{\mathcal{G}}/W_{\mathcal{H}}$. Let $\omega_{w}$ denote the $\mathcal{G}$-invariant harmonic representative of the Schubert class indexed by $w$. Then
$w\in\ \{\text{minimal rep of $W_{\mathcal{G}}/W_{\mathcal{H}}$}\}\;\;\longleftrightarrow\;\;\omega_{w}\in\{\text{Schubert harmonic basis on }\mathcal{G}/\mathcal{H}\}.$ (5.2)
The degree of $\omega_{w}$ is
@@
with $\ell(w)$ the Coxeter length. In particular, the identity coset has minimal representative $e$, corresponding to the 0-form $\omega_{e}=1$.
@@
$w\in\ \{\text{minimal rep of $W_{\mathcal{G}}/W_{\mathcal{H}}$}\}\;\;\longleftrightarrow\;\;\text{wt}_{w}=w(\mathbf{k})\in O(\mathbf{k}),$ (5.4)
@@
$\Bigl\{\;(\Omega_{p}\wedge\omega_{w_{p}})_{\text{harm}}\;\Big|\;w_{p}\in W_{H^{s}}/W_{Z_{H^{s}}(\overline{\mathbf{k}}_{i_{p}})},\;p=1,\ldots,N\Bigr\},$ (5.8)
realizing the $(H^{\vee})^{s}$-representation $V_{\Lambda_{q}}$. Here $(\Omega_{p}\wedge\omega_{w_{p}})_{\text{harm}}$ denotes the harmonic representative of the cohomology class $\Omega_{p}\wedge\omega_{w_{p}}]$.
@@
$\text{wt}((\Omega_{p}\wedge\omega_{w_{p}})_{\text{harm}})=\text{wt}(\Omega_{p}\wedge\omega_{w_{p}})=\text{wt}(\omega_{w_{p}})=w_{p}(\overline{\mathbf{k}}_{i_{p}})\in O(\overline{\mathbf{k}}_{i_{p}}).$ (5.10)
@@
$\deg(\Omega_{p})=\begin{cases}\text{dim}\mathcal{M}_{\mathrm{rel}}(q,\Phi_{0},k_{i}),&\text{if $q$ is a unite vector},\\[6.0pt] \frac{1}{2}\,\text{dim}\mathcal{M}_{\mathrm{rel}}(q,\Phi_{0},k_{i}),&\text{otherwise}.\end{cases}$ (5.11)
If $q$ is a unit vector, $\mathcal{M}_{\mathrm{rel}}(q,\Phi_{0})$ is compact, and the fiber contributes its top-degree cohomology class (the Thom class). Otherwise, $\mathcal{M}_{\mathrm{rel}}(q,\Phi_{0})$ is noncompact and the bound state on $\mathcal{M}_{\mathrm{rel}}(q,\Phi_{0},k_{i})$ is represented by the (anti-)self-dual harmonic form in middle degree [4].
@@
In what follows, for examples in Subsections 3.1 and 3.2, we explicitly compute the weights carried by the harmonic forms in the associated $(H^{\vee})^{s}$-representation. In particular, in Subsection 3.1, where $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})=C(k_{1})$ is a compact Hermitian symmetric space equipped with its standard invariant Kähler metric, all harmonic forms are parallel. Moreover, the Schubert harmonic representatives $\{\omega_{w}\}$ are orthogonal [59]. Using this orthogonal parallel harmonic basis, we construct the magnetic generators of the $(H^{\vee})^{s}$-action on the monopole Hilbert space via the matrix units $E_{m,n}$ in (4.60). When $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})$ is multi-stratified, we do not have a general proof for the parallelness of harmonic forms, so (4.60) may not apply.
@@ 5.1 $(H^{\vee})^{s}$-weight decomposition of harmonic forms
    $\overline{\mathbf{k}}_{1}=\bm{e}_{N}-\frac{1}{N}\sum^{N}_{a=1}\bm{e}_{a},$ (5.13)
@@
    Let $\omega$ be the Kähler form on $\mathbb{CP}^{N-1}$. A Schubert basis of harmonic forms is
@@
    which is a minuscule coweight of $SO(2N+1)$, and is the lowest weight of the $\mathbf{2N}$-representation in $(H^{\vee})^{s}=USp(2N)$.
@@
    A Schubert basis of harmonic forms is
@@
    $w_{-k}(\overline{\mathbf{k}}_{1})=-\bm{e}_{k},\qquad w_{+k}(\overline{\mathbf{k}}_{1})=\bm{e}_{k},$ (5.30)
@@
    $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})\cong SU(N+1)/S[U(N-1)\times U(2)]\cong\mathrm{Gr}(2,N+1),\;\;\text{dim}_{\mathbb{R}}\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})=4N-4,$ (5.35)
@@
[…]
    $f_{ij}=\sum_{m=0}^{[\frac{j-i-1}{2}]}(-1)^{m}\binom{j-i-1-m}{m}\;\omega^{\,j-i-1-2m}\wedge\Omega^{\,i-1+m}.$ (5.37)
    Minimal-length representatives of $W_{H^{s}}/W_{Z_{H^{s}}(\overline{\mathbf{k}}_{1})}$ can be chosen as
@@
    The Schubert harmonic basis may be chosen as
    $\{1,\omega,\ldots,\omega^{N-2},\omega^{N-1}+\psi,\omega^{N-1}-\psi,\omega^{N},\ldots,\omega^{2N-2}\}=\{f_{m}\}_{m=1}^{2N},$ (5.46)
    where $\omega$ is the Kähler form and $\psi$ is the unique primitive middle-degree harmonic form satisfying $\omega\wedge\psi=0$, normalized so that $\omega^{N-1}\pm\psi$ are two middle-degree Schubert harmonic forms. The associated minimal-length representatives in $W_{H^{s}}/W_{Z_{H^{s}}(\overline{\mathbf{k}}_{1})}$ are
@@
    $w_{-k}(\overline{\mathbf{k}}_{1})=-\bm{e}_{k},\qquad w_{+k}(\overline{\mathbf{k}}_{1})=\bm{e}_{k},$ (5.50)
    the weights carried by harmonic forms in the $SO(2N)$ vector representation are given by
@@
    is the lowest weight of $(\mathbf{2N{+}1})$-representation of $(H^{\vee})^{s}=SO(2N+1)$, but is not a minuscule coweight of $USp(2N)$. There is a second solution
@@
    The Euler characteristic is
@@
    Since $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})$ is homeomorphic to $\mathbb{CP}^{2N}$, they have the isomorphic cohomology rings. The basis of harmonic forms on $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})$ can be taken as [^11] Here we take $\mathbb{CP}^{2N-1}\cup\mathbb{C}^{2N}$ as $\mathbb{C}^{2N}$ with the sphere $\mathbb{S}^{4N-1}_{\infty}$ quotiented by the Hopf $\mathbb{S}^{1}$-action, so that $\mathbb{CP}^{2N-1}$ arises at infinity. A form on $\mathbb{C}^{2N}$ descends, upon restriction to $\mathbb{S}^{4N-1}_{\infty}$, to a well-defined form on $\mathbb{CP}^{2N-1}$ if and only if its restriction is basic for the Hopf fibration, i.e. $\mathbb{S}^{1}$-invariant and horizontal. Moreover, a harmonic form on $\mathbb{C}^{2N}$ whose boundary values on $\mathbb{S}^{4N-1}_{\infty}$ stay finite as $r\to\infty$ is necessarily parallel. These conditions single out (5.57), where $\tilde{\omega}=\frac{i}{2}\sum_{a=1}^{2N}dz^{a}\wedge d\bar{z}^{a}$ is the flat Kähler form on $\mathbb{C}^{2N}$.
@@
    Let $f_{k}=\tilde{\omega}^{k-1}$, which are parallel and mutually orthogonal. The simple-root generators of $SO(2N+1)$ in this basis are
@@
    As shown in (2), $C(k_{1})\cong SU(N+1)/S[U(N)\times U(1)]\cong\mathbb{CP}^{N}$, $C(k_{2})\cong SU(N+1)/S[U(N-1)\times U(2)]\cong\mathrm{Gr}(2,N+1)$,
@@
    where $\mathrm{Tot}(S^{\vee})$ is locally isomorphic to $\mathbb{C}^{2}\times\mathrm{Gr}(2,N+1)$. $\text{dim}_{\mathbb{R}}\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})=4N$.
@@
    $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})$ and $\mathrm{Gr}(2,N{+}2)$ have isomorphic cohomology rings. Let $\tilde{\omega}$] and $\tilde{\Omega}$] be the images of the degree-$2$ and degree-$4$ special Schubert generators under this isomorphism. Then the cohomology classes on $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})$ are $f^{\prime}_{ij}]$ for $1\leq i<j\leq N+2$, where
    $f^{\prime}_{ij}=\sum_{m=0}^{[\frac{j-i-1}{2}]}(-1)^{m}\binom{j-i-1-m}{m}\;\tilde{\omega}^{\,j-i-1-2m}\wedge\tilde{\Omega}^{\,i-1+m},$ (5.70)
    as in (5.37). Let $(f^{\prime}_{ij})_{\mathrm{harm}}$ be the harmonic representative of $f^{\prime}_{ij}]$. Setting $f_{ij}=f^{\prime}_{i\,j+1}$, the harmonic basis on $\mathcal{M}_{\mathrm{rel}}(1,\Phi_{0})$ is
    $\{(f_{ij})_{\mathrm{harm}}\mid 1\leq i\leq j\leq N+1\}=\{(f^{\prime}_{ij})_{\mathrm{harm}}\mid 1\leq i<j\leq N+2\},$ (5.71)